OpenStudy (goodday):

Find the product of 5×34

OpenStudy (usukidoll):



OpenStudy (usukidoll):

first multiply 4 x 5

OpenStudy (unklerhaukus):

5×34 = 5×30 + 5×4 = 5×3×10 + 5×4 =

OpenStudy (usukidoll):

whoa I wasn't taught that trick :O

OpenStudy (unklerhaukus):

(same technique, different notation)

OpenStudy (goodday):

1 1/7 is tht right

OpenStudy (usukidoll):

no.. let's try again what's 4 x 5

OpenStudy (goodday):

20

OpenStudy (usukidoll):

alright so we place the 0 down and bring the 2 up